Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Intermediate Probability

Expert replies
by Aman verma » Tue Feb 16, 2010 9:02 am
Q : Urn A contains six red and four black balls and Urn B contains four red and six black balls.One ball is drawn at random from Urn A and placed in Urn B.Then one ball is transferd at random from Urn B to Urn A .If one ball is now drawn at random from Urn A ,find the probability that it is red :

a)32/65

b32/55

c)23/55

d)56/65

e)31/65
Join the discussion
Source: — Problem Solving |

by ajith » Tue Feb 16, 2010 10:37 am
Aman verma wrote:Q : Urn A contains six red and four black balls and Urn B contains four red and six black balls.One ball is drawn at random from Urn A and placed in Urn B.Then one ball is transferd at random from Urn B to Urn A .If one ball is now drawn at random from Urn A ,find the probability that it is red :

a)32/65

b32/55

c)23/55

d)56/65

e)31/65
Case 1: One red ball is transferred to the second urn
Probabilty of that happening = 6/10 = 3/5

One red ball is transferred to the first urn from second
Probabilty of that happening = 5/11 = 5/11

One red ball is selected from urn A after this = 6/10

Probability that all of them happening together (all of the events are independent) = 3/5*5/11*3/5 = 9/55

Case 2: Red ball is first transferred, black ball is transferred back and red ball is chosen

Probability 6/10*6/11*5/10 = 9/55

Case 3: Black ball is first transferred, red ball is transferred back and red ball is chosen

Probability = 4/10*4/11*5/10 = 2/5*1/2*4/11 =4/55

Case 4: Black ball is first transferred , black ball is transferred back and red ball is chosen
4/10*7/11*6/10 =2/5*3/5*7/11= 42/5* 1/55 = 8.4/55


I am messing up calculation somewhere and getting 30.4/55 :(
Always borrow money from a pessimist, he doesn't expect to be paid back.
Join the discussion

by amittilak » Tue Feb 16, 2010 11:44 am
Ajith,
In your case 3, you have 4/10*4/11*5/10
I think it should be, 4/10*4/11*7/10, since you are adding 1 extra red ball from Urn B to urn A.
That gives us,
(6*5*6+6*6*5+4*4*7+4*7*6)/(10*11*10 =
27/55

What's the OA on this?
ajith wrote:
Aman verma wrote:Q : Urn A contains six red and four black balls and Urn B contains four red and six black balls.One ball is drawn at random from Urn A and placed in Urn B.Then one ball is transferd at random from Urn B to Urn A .If one ball is now drawn at random from Urn A ,find the probability that it is red :

a)32/65

b32/55

c)23/55

d)56/65

e)31/65
Case 1: One red ball is transferred to the second urn
Probabilty of that happening = 6/10 = 3/5

One red ball is transferred to the first urn from second
Probabilty of that happening = 5/11 = 5/11

One red ball is selected from urn A after this = 6/10

Probability that all of them happening together (all of the events are independent) = 3/5*5/11*3/5 = 9/55

Case 2: Red ball is first transferred, black ball is transferred back and red ball is chosen

Probability 6/10*6/11*5/10 = 9/55

Case 3: Black ball is first transferred, red ball is transferred back and red ball is chosen

Probability = 4/10*4/11*5/10 = 2/5*1/2*4/11 =4/55


Case 4: Black ball is first transferred , black ball is transferred back and red ball is chosen
4/10*7/11*6/10 =2/5*3/5*7/11= 42/5* 1/55 = 8.4/55


I am messing up calculation somewhere and getting 30.4/55 :(
Join the discussion

by jeffedwards » Tue Feb 16, 2010 12:33 pm
Wow, used the same method as the two of you above. Is there an easier method? This took me much longer that 2 minutes to solve.
Join the discussion

by Stuart@KaplanGMAT » Tue Feb 16, 2010 3:57 pm
Some minor math errors in there:

Case 1: Red moved, then red moved, then red chosen:

3/5 * 5/11 * 3/5 = 9/55

Case 2: Red moved, then black moved, then red chosen:

3/5 * 6/11 * 1/2 = 9/55

Case 3: Black moved, then red moved, then red chosen:

2/5 * 4/11 * 7/10 = 28/275

Case 4: Black moved, then black moved, then red chosen:

2/5 * 7/11 * 3/5 = 42/275

Add them up: 18/55 + 70/275 = 18/55 + 14/55 = 32/55... choose B.

* * *

I actually arrived at B by logic/estimation as well:

We're moving a ball away then a ball back; since there are more red balls in the first urn, there's a slightly better chance of decreasing the number of red balls in Urn A than increasing it.

Our original probability of drawing a red ball is 60%, so the final answer should be a bit less than 60%.

A, C and E are all under 50% - seem too low.
D is more than 60% - definitely too high.
B is just under 60% - just right!

Here's another way to use logic/estimation:

we know that our 3 denominators are 5, 11 and 5. Therefore, the denominator in the final answer must be a factor of 275.

Eliminate A, D and E.

(If nothing else, this is a super quick way to give ourselves a 50/50 shot.)

C goes from 60% original probability to just over 40% new probability - that's way too low, eliminate C.

Pick B!
Image

Stuart Kovinsky | Kaplan GMAT Faculty | Toronto

Kaplan Exclusive: The Official Test Day Experience | Ready to Take a Free Practice Test? | Kaplan/Beat the GMAT Member Discount
BTG100 for $100 off a full course
Join the discussion

by jeffedwards » Tue Feb 16, 2010 5:17 pm
Genius Stuart, I love all of your posts!
Join the discussion

by thephoenix » Tue Feb 16, 2010 9:00 pm
there are 4 cases

1)RR [red rmoved from A, red removed fromB]
2)RB
3)BR
4)BB


for 1) probability pickin a red ball from A finally=6/10 (for picking red from urn )*5/11(for picking red from urn b After the addition of a red ball from A)*6/10=9/55
similarly for 2)P=6/10 * 6/11 * 5/10 = 9/55

for 3)=28/275

for 4)=42/275

tot=32/55
Join the discussion

by sumanr84 » Wed Feb 17, 2010 12:29 am
Stuart Kovinsky wrote: (If nothing else, this is a super quick way to give ourselves a 50/50 shot.)

C goes from 60% original probability to just over 40% new probability - that's way too low, eliminate C.

Pick B!
I love this one !!
I am on a break !!
Join the discussion

by Aman verma » Wed Feb 17, 2010 2:22 am
B it is.
Join the discussion